# Models of Set Theory: Extensions and Dead-ends

### v5

@@ 6. ADDENDUM AND CORRIGENDUM TO [En-3]
(a) $V_{n}$ and $W_{n}$ remain inseparable in $\mathcal{N}$ for all $n\in\mathbb{\omega}$.
@@
(b) For each $n\in\mathbb{\omega}$ there is some $d_{n}\in N$ such that $\mathcal{N}\models d_{n}\in\mathbb{D}_{n}$ (i.e., $d_{n}$ satisfies the formula that defines $\mathbb{D}_{n}$ in $\mathcal{M}$), and for each $d\in\mathbb{D}_{n},$ $\mathcal{N}\models\left(d<_{\mathbb{D}_{n}}d_{n}\right).$
@@
Let $\mathcal{L}^{+}$ be the language $\mathcal{L}$ augmented with constants $\left\{d_{n}:n\in\mathbb{\omega}\right\}\cup\{m:m\in M\}$, and consider the $\mathcal{L}^{+}$-theory $T$ below:
$T:=\mathrm{Th}(\mathcal{M},m)_{m\in M}+\left\{d_{n}\in\mathbb{D}_{n}:n\in\mathbb{\omega}\right\}\cup\left\{d<_{\mathbb{D}_{n}}d_{n}:n\in\mathbb{\omega},\text{\ }d\in\mathbb{D}_{n}\right\}.$
@@
$T$ is readily seen to be consistent since it is finitely satisfiable in $\mathcal{M}\mathfrak{.}$ Moreover, it is clear that if $\mathcal{N}\models T$, then $\mathcal{M}\prec\mathcal{N}$ and $\mathcal{N}$ satisfies condition (b) of the theorem. To arrange a model of $T$ in which conditions (a) also holds requires a delicate omitting types argument. First, we need a pair of preliminary lemmas. The proof of Lemma A.3 is routine and left to the reader, and Lemma A.4 is an immediate corollary of Lemma A.3.
@@
A.3. Lemma. The following two conditions are equivalent for a sentence $\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$ of $\mathcal{L}^{+}$.
$(i)$   $T\vdash\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$.
@@
$(ii)$ $\ \ \mathcal{M}\models\exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{b}\in D_{k_{n}}\forall s\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{\mathbb{D}_{k_{n}}}s_{i}\right)\rightarrow\varphi(s_{1},\cdot\cdot\cdot,s_{n})\right].$
@@
A.4. Lemma. The following two conditions are equivalent for a sentence $\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$ of $\mathcal{L}^{+}$.
$(i)$   $T+\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$ is consistent.
@@
$(ii)$ $\ \ \mathcal{M}\models\forall r_{1}\in\mathbb{D}_{k_{1}}\exists s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\forall r_{b}\in\mathbb{D}_{k_{n}}\exists s\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{\mathbb{D}_{k_{n}}}s_{i}\right)\wedge\varphi(s_{1},\cdot\cdot\cdot,s_{n})\right].$
@@
We are now ready to carry out an omitting types argument to complete the proof of Lemma A.2. For each formula $\psi(y,\overline{x})$ of $\mathcal{L}^{+}$, and each $n\in\omega$, consider the following $p$-type formulated in the language $\mathcal{L}^{+}$, where $\overline{x}=\left\langle x_{1},...,x_{p}\right\rangle$ is a $p$-tuple of variables:
$\Sigma_{n}^{\psi}(\overline{x}):=\{\psi(a,\overline{x}):a\in V_{n}\}\cup\{\lnot\psi(b,\overline{x}):b\in W_{n}\}.$
Note that $\Sigma_{n}^{\psi}$ expresses that $\psi(y,\overline{x})$ separates $V_{n}$ and $W_{n}.$
@@
A.5. Lemma. $\Sigma_{n}^{\psi}$ is locally omitted by $T$ for each formula $\psi(y,\overline{x})$, and each $n\in\mathbb{\omega}$.
@@
Proof. Suppose to the contrary that there is a formula $\theta(\overline{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$ of $\mathcal{L}^{+}$ and some $n\in\omega$ such that (1) through (3) below hold:
(1) $T+\exists\overline{x}\ \theta(\overline{x},c)$ is consistent.
(2) For all $a\in V_{n},$ $T\vdash\forall\overline{x}\ \left[\theta(\overline{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})\rightarrow\psi(a,\overline{x})\right].$
@@
(3) For all $b\in W_{n},$ $T\vdash\forall\overline{x}\ \left[\theta(\overline{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})\rightarrow\lnot\psi(b,\overline{x})\right].$
@@
$(1^{\prime})$ $\ \mathcal{M}\models\forall r_{1}\in\mathbb{D}_{k_{1}}\exists s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\forall r_{b}\in\mathbb{D}_{k_{n}}\exists s_{n}\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{D_{n}}s_{i}\right)\wedge\exists\overline{x}\ \theta(\overline{x},\overline{s})\right],$ where $\overline{s}=\left\langle s_{1},...,s_{n}\right\rangle.$
$(2^{\prime})$ For all $a\in V_{n}$, $\mathcal{M}\models\lambda(a),$ where:
$\lambda(\mathbf{v}):=\forall\overline{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{b}\in\mathbb{D}_{k_{n}}\forall s_{n}\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{D_{n}}s_{i}\right)\rightarrow\left(\theta(\overline{x},\overline{s})\rightarrow\psi(\mathbf{v},\overline{x})\right)\right].$
$(3^{\prime})$ For all $b\in W_{n},\ \mathcal{M}\models\gamma(b)$, where:
$\gamma(\mathbf{w}):=\forall\overline{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{b}\in\mathbb{D}_{k_{n}}\forall s_{n}\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{D_{n}}s_{i}\right)\rightarrow\left(\theta(\overline{x},\overline{s})\rightarrow\lnot\psi(\mathbf{w},\overline{x})\right)\right].$
@@
(4) $\mathcal{M}\models\lambda(m)\wedge\gamma(m).$
It is easy to see, using the fact that each $\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right)$ is a directed that (4) implies:
(5)))) $\mathcal{M}\models\exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{b}\in\mathbb{D}_{k_{n}}\forall s\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{\mathbb{D}_{n}}s_{i}\right)\rightarrow\left(\psi(m,\overline{x})\wedge\lnot\psi(m,\overline{x})\right)\right].$
Since each $\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right)$ is a directed set with no last element, (5) yields a contradiction, so the proof is complete. $\square$ (Lemma A.5)
@@
In this stage we employ the result of the first stage together with some set-theoretical considerations to establish the Rubin-Shelah-Schmerl Theorem. The main idea here is a variant of the one used by Schmerl, [Schm-2], which itself based on a method of $\diamondsuit_{\mathbb{\aleph}_{1}}$-elimination introduced by Shelah [Sh]. This stage has three steps.
Step 1 of Stage 2. It is well-known that there is an $\omega_{1}$-closed notion $\mathbb{Q}_{1}$ in the universe $\mathrm{V}$ of $\mathrm{ZFC}$ such that the $\mathbb{Q}_{1}$-generic extension $\mathrm{V}^{\mathbb{Q}_{1}}$ of $\mathrm{V}$ satisfies $\mathrm{ZFC}+\diamondsuit_{\mathbb{\aleph}_{1}}$ [Ku, Theorem 8.3] (note that $\aleph_{1}$ is absolute between $\mathrm{V}$ and $\mathrm{V}^{\mathbb{Q}_{1}}$ since $\mathbb{Q}_{1}$ is $\omega_{1}$-closed). Thus by Step 1, given a model $\mathcal{M}$ in the universe $\mathrm{V}$ of $\mathrm{ZF}$, the forcing extension $\mathrm{V}^{\mathbb{Q}_{1}}$ satisfies “there is an elementary extension $\mathcal{N}$ of $\mathcal{M}$ such that $\mathcal{N}$ is a Rubin model”. In $\mathrm{V}^{\mathbb{Q}_{1}}$, since $\mathcal{N}$ is a Rubin model, for each $s\in N$ that is infinite in the sense of $\mathcal{N}$, there is a subset $C_{s}=\left\{c_{\alpha}^{s}:\alpha\in\omega_{1}\right\}$ of $\mathcal{N}$ such that $\left\langle c_{\alpha}^{s}:\alpha\in\omega_{1}\right\rangle$ is an $\omega_{1}$-chain that is cofinal in the directed set $\left([s]^{<\omega}\right)^{\mathcal{M}}.$
@@
Note that the $\mathcal{N}^{+}$-definable ranked trees include trees of the form $\tau_{s}$, as well as the ranked tree $\tau_{\mathrm{class}}^{\mathcal{N}}$ (whose branches uniquely determine piecewise coded subsets of $\mathcal{N}$, as noted in Remark 5.14).
@@
Proof. Suppose $B$ is a branch of $\tau$. Then since ($\mathbb{L},\ \leq_{\mathbb{L}})$ has cofinality $\omega_{1},$ there is a subset $B_{0}$ of $B$ of order type $\omega_{1}$ that is cofinal in $B$. Therefore there is some unbounded $S\subseteq\omega_{1}$ such that $f$ is constant on $\left\{f(b):b\in B_{0}\right\}\}.$ Fix some element $b_{0}\in B_{0}$ and consider subset $X$ of $\mathbb{T}$ that is defined in $\left(\tau,f\right)$ by the formula:
@@
By the defining property of weakly specializing functions, $X$ is linearly ordered and is cofinal in $B$. This makes it clear that $B$ is definable in $\left(\tau,f\right)$ as the downward closure of $X$, i.e., by the formula $\psi(x,b_{0}):=\left[\exists y(\varphi(y,b_{0})\wedge x\leq_{\mathbb{T}}y\right].$square
@@
By a remarkable theorem, due independently to Baumgartner [Bau] and Shelah [Sh], given any ranked tree $\tau$ of cardinality $\aleph_{1}$ and of cofinality $\omega_{1}$, there is a c.c.c. notion of forcing $\mathbb{Q}_{\tau}$ such that $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}_{\tau}}:=\left(\mathrm{V}^{\mathbb{Q}_{1}}\right)^{\mathbb{Q}_{\tau}}$ contains a function $f_{\tau}$ such that weakly specializes $\tau.$ [^24] This result generalizes the fact, established by Baumgartner, Malitz, and Reinhardt [BMR] that every $\omega_{1}$-Aronszajn tree can be turned into a special Aronszajn tree in a c.c.c. forcing extension of the universe. The forcing condition in $\mathbb{Q}_{\tau}$ consists of all finite attempts to build a function that weakly specializes $\tau$. Moreover, given any collection $\{\tau_{\alpha}:\alpha\in\kappa\}$ of ranked trees of cofinality $\omega_{1}$, the finite support product $\mathbb{Q}$ of $\left\{\mathbb{Q}_{\tau_{\alpha}}:\alpha\in\kappa\right\}$ is also known to have the c.c.c. property [^25] It is well-known that the finite support product of a given collection of forcing notions $\left\{\mathbb{P}_{\alpha}:\alpha\in\kappa\right\}$ has the c.c.c. property iff every finite sub-product of $\left\{\mathbb{P}_{\alpha}:\alpha\in\kappa\right\}$ has the c.c.c. property., as indicated in [CZ, Corollary 3.3]262626[CZ, Corollary 3.3] is overtly about specializing $\omega_{1}$-Aronszajn trees, but as shown by Baumgartner [Bau, Section 7], given a tree $\tau$ of height $\omega_{1}$ with at most $\aleph_{1}$ many branches, there is an Aronszajn subtree $\tau_{0}$ of $\tau$ of height $\omega_{1}$, and moreover, any function $f_{0}$ that specializes $\tau_{0}$ can be canonically extended to a function $f$ that specializes $\tau.$ For more detail, see the proof of [Sw, Lemma 3.5].. Note that for each $\alpha\in\kappa$, $\mathbb{Q}$ forces the existence of a function $f_{\tau_{\alpha}}$ that weakly specializes $\tau_{\alpha}.$
@@
  • Thus by choosing $\{\tau_{\alpha}:\alpha\in\omega_{1}\}$ to be the collection of ranked trees that are parametrically definable in $\mathcal{N}^{+}$, there is a function $g(s,x)$ in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}:=\left(\mathrm{V}^{\mathbb{Q}_{1}}\right)^{\mathbb{Q}}$ such that for each infinite set $s$ of $\mathcal{N}$, $g(s,x)$ weakly specializes $\tau_{s}$ ($\tau_{s}$ was defined in Step 1 of Stage 2), and a function $h$ that weakly specializes $\tau_{\mathrm{class}}^{\mathcal{N}}.$
@@
The next lemma shows that forcing with $\mathbb{Q}$ does not add new branches to ranked trees in the ground model that have cofinality $\omega_{1}$. Note that when the lemma below is applied to the case when V is replaced by $\mathrm{V}^{\mathbb{Q}_{1}}$, it shows that the ample supply of rather branchless models in $\mathrm{V}^{\mathbb{Q}_{1}}$ remain rather branchless after forcing with $\mathbb{Q}$, i.e., in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}.$

### v6

@@ 6. ADDENDUM AND CORRIGENDUM TO [En-3]
  1. (a)
    $V_{n}$ and $W_{n}$ remain inseparable in $\mathcal{N}$ for all $n\in\mathbb{\omega}$.
  2. (b)
@@
    For each $n\in\mathbb{\omega}$ there is some $d_{n}\in N$ such that $\mathcal{N}\models d_{n}\in\mathbb{D}_{n}$ (i.e., $d_{n}$ satisfies the formula that defines $\mathbb{D}_{n}$ in $\mathcal{M}$), and for each $m\in\mathbb{D}_{n},$ $\mathcal{N}\models\left(m<_{\mathbb{D}_{n}}d_{n}\right).$
@@
Let $\mathcal{L}^{+}$ be the language $\mathcal{L}$ augmented with constants from $C_{1}=\{\dot{m}:m\in M\}$ and $C_{2}=\left\{d_{n}:n\in\mathbb{\omega}\right\}$ (where $C_{1}$ and $C_{2}$ are disjoint), and consider the $\mathcal{L}^{+}$-theory $T$ below:
$T:=\mathrm{Th}(\mathcal{M},m)_{m\in M}+\left\{d_{n}\in\mathbb{D}_{n}:n\in\mathbb{\omega}\right\}\cup\left\{\dot{m}<_{\mathbb{D}_{n}}d_{n}:\mathcal{N}\models\left(m\in\mathbb{D}_{n}\right),n\in\mathbb{\omega}\right\}.$
@@
$T$ is readily seen to be consistent since it is finitely satisfiable in $\mathcal{M}\mathfrak{.}$ Moreover, it is clear that if $\mathcal{N}\models T$, then $\mathcal{M}\prec\mathcal{N}$ and $\mathcal{N}$ satisfies condition (b) of the theorem. To arrange a model of $T$ in which conditions (a) also holds requires a delicate omitting types argument. First, we need a pair of preliminary lemmas. The proofs of Lemma A.3 and A.4 are routine and left to the reader. Note that they are each other’s contrapositive, so it is sufficient to only verify one of them.
@@
A.3. Lemma. The following two conditions are equivalent for a sentence $\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ of $\mathcal{L}^{+}$.
  1. $(i)$
    $T\vdash\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$.
  2. $(ii)$
@@
    $\mathcal{M}\models\exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{p}\in\mathbb{D}_{k_{p}}\forall s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\rightarrow\varphi(s_{1},\cdot\cdot\cdot,s_{p})\right].$
@@
A.4. Lemma. The following two conditions are equivalent for a sentence $\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ of $\mathcal{L}^{+}$.
  1. $(i)$
    $T+\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ is consistent.
  2. $(ii)$
@@
    $\mathcal{M}\models\forall r_{1}\in\mathbb{D}_{k_{1}}\exists s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\forall r_{p}\in\mathbb{D}_{k_{p}}\exists s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\wedge\varphi(s_{1},\cdot\cdot\cdot,s_{p})\right].$
@@
We are now ready to carry out an omitting types argument to complete the proof of Lemma A.2. For each formula $\psi(y,\vec{x})$ of $\mathcal{L}^{+}$, and each $n\in\omega$, consider the following $q$-type formulated in the language $\mathcal{L}^{+}$, where $\vec{x}=\left\langle x_{1},\cdot\cdot\cdot,x_{q}\right\rangle$ is a $q$-tuple of variables:
$\Sigma_{n}^{\psi}(\vec{x}):=\{\psi(\dot{a},\vec{x}):a\in V_{n}\}\cup\{\lnot\psi(\dot{b},\vec{x}):b\in W_{n}\}.$
Note that $\Sigma_{n}^{\psi}$ expresses that $\psi(y,\vec{x})$ separates $V_{n}$ and $W_{n}.$
@@
A.5. Lemma. $\Sigma_{n}^{\psi}$ is locally omitted by $T$ for each formula $\psi(y,\vec{x})$, and each $n\in\mathbb{\omega}$.
@@
Proof. Suppose to the contrary that there is a formula $\theta(\vec{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ of $\mathcal{L}^{+}$ and some $n\in\omega$ such that (1) through (3) below hold:
  1. $(1)$
    $T+\exists\vec{x}\ \theta(\vec{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ is consistent.
  2. $(2)$
    For all $a\in V_{n},$ $T\vdash\forall\vec{x}\left[\theta(\vec{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})\rightarrow\psi(\dot{a},\vec{x})\right].$
  3. $(3)$
@@
    For all $b\in W_{n},$ $T\vdash\forall\vec{x}\left[\theta(\vec{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})\rightarrow\lnot\psi(\dot{b},\vec{x})\right].$
@@
  1. $(1^{\prime})$
    $\mathcal{M}\models\forall r_{1}\in\mathbb{D}_{k_{1}}\exists s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\forall r_{p}\in\mathbb{D}_{k_{p}}\exists s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\wedge\exists\vec{x}\ \theta(\vec{x},\vec{s})\right],$ where $\vec{s}=\left\langle s_{1},\cdot\cdot\cdot,s_{p}\right\rangle.$
  2. $(2^{\prime})$
    For all $a\in V_{n}$, $\mathcal{M}\models\lambda(a),$ where:
    $\lambda(\mathbf{v}):=\forall\vec{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{p}\in\mathbb{D}_{k_{p}}\forall s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\rightarrow\left(\theta(\vec{x},\vec{s})\rightarrow\psi(\mathbf{v},\vec{x}\ )\right)\right].$
  3. $(3^{\prime})$
    For all $b\in W_{n},\ \mathcal{M}\models\gamma(b)$, where:
    $\gamma(\mathbf{w}):=\forall\vec{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{n}\in\mathbb{D}_{k_{p}}\forall s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\rightarrow\left(\theta(\vec{x},\vec{s})\rightarrow\lnot\psi(\mathbf{w},\vec{x})\right)\right].$
@@ 3. $(3^{\prime})$
  1. $(4)$
    $\mathcal{M}\models\lambda(m)\wedge\gamma(m).$
Since each $\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right)$ is a directed set, (4) implies:
  1. $(5)$
    $\mathcal{M}\models\forall\vec{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{p}\in\mathbb{D}_{k_{p}}\forall s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\rightarrow\left(\psi(m,\vec{x})\wedge\lnot\psi(m,\vec{x})\right)\right].$
Recall that no $\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right)$ has a last element. This shows that (5) yields a contradiction, so the proof is complete.
$\square$ (Lemma A.5)
@@ 1. $(5)$
In this stage we employ the result of the first stage together with some set-theoretical considerations to establish the Rubin-Shelah-Schmerl Theorem. The main idea here is a variant of the one used by Schmerl, [Schm-2], which itself based on a method of $\diamondsuit_{\omega_{1}}$-elimination introduced by Shelah [Sh]. This stage has three steps.
Step 1 of Stage 2. It is well-known that there is an $\omega_{1}$-closed notion $\mathbb{Q}_{1}$ in the universe $\mathrm{V}$ of $\mathsf{ZFC}$ such that the $\mathbb{Q}_{1}$-generic extension $\mathrm{V}^{\mathbb{Q}_{1}}$ of $\mathrm{V}$ satisfies $\mathsf{ZFC}+\diamondsuit_{\mathbb{\omega}_{1}}$ [Ku, Theorem 8.3] (note that $\aleph_{1}$ is absolute between $\mathrm{V}$ and $\mathrm{V}^{\mathbb{Q}_{1}}$ since $\mathbb{Q}_{1}$ is $\omega_{1}$-closed). Thus by Step 1, given a model $\mathcal{M}$ in the universe $\mathrm{V}$ of $\mathsf{ZF}$, the forcing extension $\mathrm{V}^{\mathbb{Q}_{1}}$ satisfies “there is an elementary extension $\mathcal{N}$ of $\mathcal{M}$ such that $\mathcal{N}$ is a Rubin model”. In $\mathrm{V}^{\mathbb{Q}_{1}}$, since $\mathcal{N}$ is a Rubin model, for each $s\in N$ that is infinite in the sense of $\mathcal{N}$, there is a subset $C_{s}=\left\{c_{\alpha}^{s}:\alpha\in\omega_{1}\right\}$ of $\mathcal{N}$ such that $\left\langle c_{\alpha}^{s}:\alpha\in\omega_{1}\right\rangle$ is an $\omega_{1}$-chain that is cofinal in the directed set $\left([s]^{<\omega}\right)^{\mathcal{M}}.$
@@
Note that the $\mathcal{N}^{+}$-definable ranked trees include trees of the form $\tau_{s}$, as well as the ranked tree $\tau_{\mathrm{class}}^{\mathcal{N}}$ (whose branches uniquely determine classes of $\mathcal{N}$, as noted in Remark 5.14).
@@
Proof. Suppose $B$ is a branch of $\tau$. Then since ($\mathbb{L},\ \leq_{\mathbb{L}})$ has cofinality $\omega_{1},$ there is a subset $S$ of $B$ of order type $\omega_{1}$ that is unbounded in $B$. Therefore there is some subset $B_{0}\$of $B$ that is unbounded in $B$ such that $f$ is constant on $B_{0}.$ Fix some element $b_{0}\in B_{0}$ and consider subset $X$ of $\mathbb{T}$ that is defined in $\left(\tau,f\right)$ by the formula:
@@
By the defining property of weakly specializing functions, $X$ is linearly ordered and is unbounded in $B$. This makes it clear that $B$ is definable in $\left(\tau,f\right)$ as the downward closure of $X$, i.e., by the formula $\psi(x,b_{0}):=\exists y\left[\varphi(y,b_{0})\wedge x\leq_{\mathbb{T}}y\right].$ $\square$
@@
By a remarkable theorem, due independently to Baumgartner [Bau] and Shelah [Sh], given any ranked tree $\tau$ of cardinality $\aleph_{1}$ and of cofinality $\omega_{1}$, there is a c.c.c. notion of forcing $\mathbb{Q}_{\tau}$ such that $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}_{\tau}}:=\left(\mathrm{V}^{\mathbb{Q}_{1}}\right)^{\mathbb{Q}_{\tau}}$ contains a function $f_{\tau}$ such that weakly specializes $\tau.$ [^37] This result generalizes the fact, established by Baumgartner, Malitz, and Reinhardt [BMR] that every $\omega_{1}$-Aronszajn tree can be turned into a special Aronszajn tree in a c.c.c. forcing extension of the universe. Here $\mathbb{Q}_{\tau}$ consists of all finite attempts for building a function that weakly specializes $\tau$ (ordered by inclusion). Moreover, given any collection $\{\tau_{\alpha}:\alpha\in\kappa\}$ of ranked trees of cofinality $\omega_{1}$, the finite support product $\mathbb{Q}$ of $\left\{\mathbb{Q}_{\tau_{\alpha}}:\alpha\in\kappa\right\}$ is also known to have the c.c.c. property [^38] It is well-known that the finite support product of a given collection of forcing notions $\left\{\mathbb{P}_{\alpha}:\alpha\in\kappa\right\}$ has the c.c.c. property iff every finite sub-product of $\left\{\mathbb{P}_{\alpha}:\alpha\in\kappa\right\}$ has the c.c.c. property., as indicated in [CZ, Corollary 3.3]393939[CZ, Corollary 3.3] is overtly about specializing $\omega_{1}$-Aronszajn trees, but as shown by Baumgartner [Bau, Section 7], given a tree $\tau$ of height $\omega_{1}$ with at most $\aleph_{1}$ many branches, there is an Aronszajn subtree $\tau_{0}$ of $\tau$ of height $\omega_{1}$, and moreover, any function $f_{0}$ that specializes $\tau_{0}$ can be canonically extended to a function $f$ that specializes $\tau.$ For more detail, see the proof of [Sw, Lemma 3.5].. Note that for each $\alpha\in\kappa$, $\mathbb{Q}$ forces the existence of a function $f_{\tau_{\alpha}}$ that weakly specializes $\tau_{\alpha}.$
@@
  • Thus by choosing $\{\tau_{\alpha}:\alpha\in\omega_{1}\}$ to be the collection of ranked trees that are parametrically definable in $\mathcal{N}^{+}$, there is a function $g(s,x)$ in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}:=\left(\mathrm{V}^{\mathbb{Q}_{1}}\right)^{\mathbb{Q}}$ such that for each infinite set $s$ of $\mathcal{N}$, $g(s,x)$ weakly specializes $\tau_{s}$ ($\tau_{s}$ was defined in Step 1 of Stage 2), and a function $f$ that weakly specializes $\tau_{\mathrm{class}}^{\mathcal{N}}.$
@@
The next lemma shows that forcing with $\mathbb{Q}$ does not add new branches to ranked trees in the ground model that have cofinality $\omega_{1}$. Note that when the lemma below is applied to the case when V is replaced by $\mathrm{V}^{\mathbb{Q}_{1}}$
, it shows that the ample supply of rather branchless models in $\mathrm{V}^{\mathbb{Q}_{1}}$ remain rather branchless after forcing with $\mathbb{Q}$, i.e., in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}.$